OpenStudy (anonymous):

an enigneer worked for 6 days and her assistant worked for 7 days on a project. Together they received a slary of $4500.00. The next week the engineer worked 5 days and the assistant 3 days and earned a combined salary of $2900.00. What is the daily salary for each? Engineer $______ Assistant $ _____

OpenStudy (anonymous):

Let X be equal to the daily salary of the engineer Let Y be equal to the daily salary of the assistant for the first week, the engineer worked 6 days and the assistant worked 7 days to make a total of 4500 this means that 6X+7Y=4500 do you understand where i got this equation from?

OpenStudy (anonymous):

6x= 6 days for the engineer. 7y= 7 days for assistant=4500 received salary

OpenStudy (anonymous):

ok how about for the 2nd week?

OpenStudy (anonymous):

3x+5y=2900.00

OpenStudy (anonymous):

x is the engineer salary, y is the assistant the engineer worked 5 days and the assistant worked only 3

OpenStudy (anonymous):

my bad yeah i was looking at the wrong thing

OpenStudy (anonymous):

now you have 2 equations with 2 unknowns can you solve it now?

OpenStudy (anonymous):

umm imma need some guidence

OpenStudy (anonymous):

6x+7y=4500 5x+3y=2900 ignore the decimal places for now take the first equation and solve for x the substitute x back in for the 2nd equation and then solve for y once you have y, substitute that back into the first or 2nd equation to determine x

OpenStudy (anonymous):

there are many other ways of solving it but substitution seems easier to explain

OpenStudy (anonymous):

okay once i do that ill come back to you

OpenStudy (anonymous):

how do you solve for x? like where am i supposed to start from?

OpenStudy (anonymous):

ug, why do they have to be such ugly numbers 6x+7y=4500 7y=4500-6x y=(4500-6x)/7

OpenStudy (anonymous):

or 6x+7y=4500 6x=4500-7y x=(4500-7y)/6

OpenStudy (anonymous):

plug either one of those into the 2nd equation and solve for the remaining variable

OpenStudy (anonymous):

okay ill get back to you when i get stuck

OpenStudy (anonymous):

okay so i choose to do 6x+7y=4500 6x=4500-7y x=(4500-7y)/6. so what do you do? like just pop that last part in the calculator? or figure it out?

OpenStudy (anonymous):

then substitute it into the 2nd equation 5((4500-7y)/6)+3y=2900 simplify, and solve for y

OpenStudy (anonymous):

okay that seems like alot how am i supposed to do that?

OpenStudy (anonymous):

you distribute and use a calculator in this case (5*4500/6)-(5*7/6)y +3y = 2900

OpenStudy (anonymous):

so pop that in the calculator without the "=2900"

OpenStudy (anonymous):

what is that y goanna go for? can i put it as 1?

OpenStudy (anonymous):

you cant stick variables into a calculator.....

OpenStudy (anonymous):

i knwo that. thats why im asking can i put them in as 1?

OpenStudy (anonymous):

(5*4500/6)-(5*7/6)y +3y = 2900 -(5*7/6)y +3y = 2900-(5*4500/6) SIMPLIFY INTO DECIMALS then solve for y

OpenStudy (anonymous):

you are trying to determine y so no you cannot just put a random value for it

OpenStudy (anonymous):

im sorry but i just dont undersdtand. :/

OpenStudy (anonymous):

-5*7/6=???

OpenStudy (anonymous):

-5.8

OpenStudy (anonymous):

-5.8+3=?

OpenStudy (anonymous):

2900-(5*4500/6)=???

OpenStudy (anonymous):

-2.8 and -850

OpenStudy (anonymous):

-2.8y=-850

OpenStudy (anonymous):

y=?????

OpenStudy (anonymous):

ohhhhh didvide sorry

OpenStudy (anonymous):

umm its -304.67?????

OpenStudy (anonymous):

it should be positive anyways now you have your y value go solve for the x value 6x+7y=4500

OpenStudy (anonymous):

so the y wowuld be for the assistant?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay now to figure out the engineer i have to do the same thing?

OpenStudy (anonymous):

no, since you already have a value for y you substitute it back into one of the original equations 6x+7(304.67)=4500 then solve for x you could do the same thing for the engineer, but you would end up doing more work than you would have to

OpenStudy (anonymous):

ughh i always get confused when its time to put in the calculator im sorry but with the 6x again what do i do?

OpenStudy (anonymous):

4500-7*304.67

OpenStudy (anonymous):

2367.31

OpenStudy (anonymous):

6x=2367.31

OpenStudy (anonymous):

should i round?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

405.65??

OpenStudy (anonymous):

thats your answer for the engineer

OpenStudy (anonymous):

yes....

OpenStudy (anonymous):

is it right?

OpenStudy (anonymous):

no idea, didnt actually do the problem

OpenStudy (anonymous):

i dont think it is